please help me please..............
if (x+7)(x-3)=0,then x=? options are
(A)7 OR 3
(B)7 OR -3
(C)-7 OR 3
(D)-7 OR -3
(E) -4 OR -3 PLEASE TELL ME ACCORDING TO THE OPTIONS PLEASE

Answer by ewatrrr(24785)

(x+7)(x-3)=0
(x+7) = 0 , x = -7
0r (x-3)=0 , x = 3
Option(C)